OpenStudy (anonymous):

find the fixed points or points of the function if any g(x)=x^2-4x+1

OpenStudy (anonymous):

what do you mean fixed points?

OpenStudy (anonymous):

i think he meant roots

OpenStudy (anonymous):

no it says fixed points of a function

OpenStudy (anonymous):

what is fixed points of a functin..?

OpenStudy (anonymous):

A fixed point is a point that's mapped to itself.

OpenStudy (anonymous):

But I don't know if they have root or not that is why I am asking

OpenStudy (anonymous):

so how do you find that?

OpenStudy (anonymous):

we cant find its roots so we use b^2-4ac= delta

OpenStudy (anonymous):

then x should be equal y \[x=x^2-4x+1\] \[x^2-5x+1=0\] I hope you can solve this quadratic equation yourself and answer will be \[(\frac{5-\sqrt{21}}{2};\frac{5-\sqrt{21}}{2})\] \[(\frac{5+\sqrt{21}}{2};\frac{5+\sqrt{21}}{2})\]

OpenStudy (anonymous):

you complete me guy

OpenStudy (anonymous):

I will try thanks

OpenStudy (anonymous):

I had help with this earlier but I had a hard time following.